OpenStudy (anonymous):

-2x-5y=-13 find the value of y when x equals -1

OpenStudy (anonymous):

-2(-1)-5y=-13 2-5y=-13 -5y=-15 y=3

OpenStudy (anonymous):

-2x - 5y = -13 (when x = -1) so now you sub -1 in for x -2(-1) - 5y = -13 2 - 5y = -13 -5y = -13 - 2 -5y = -15 u = 3

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

:)